Average age of 30 students in a class is 10. then 1 more is took admission in that class. now the average of their age is 11. then find the age of last one?

Answer by balthazar612(14)

To solve this we know that,
Average of age= total ages of students/ total Students
10 = total ages of students/30
10*30= total ages of students
300= Total ages of students

Now
When 1 student is added there will be total 11 students using above formula
Average of age= Total ages of students/ total ages of students
11 = total ages of students/ 31
Total ages of students= 31*11
= 341
Change in total age= 341 - 300

Age of 11th student= 41
Answer= 41